A particle has a velocity of v=(-2.0i +3.5j)m/s. The particle starts at r= ( 1.5i -3.1j)m at t=0. Give the position and acceleration as a function of time. What is the shape of the resulting path?


1
Expert's answer
2021-01-19T07:08:41-0500

(1)


"a=\frac{dv}{dt}=\frac{d(-2i+3.5j)}{dt}=0(m/s^2)" . Answer


(2)


"v=\frac{dr}{dt}\to dr=vdt\to r(t)=\int vdt=\int{(-2i+3.5j)dt}="


"=-2ti+3.5tj+r_0=-2ti+3.5tj+1.5i-3.1j="


"=(-2t+1.5)i+(3.5t-3.1)j" . Answer


"x=-2t+1.5" and "y=3.5t-3.1"


We get


"y=-1.75x-0.475" - straight line. Answer